OpenStudy (anonymous):

Let f(x) = 3x +2 and g(x) = 7x + 6. Find f • g and its domain.

OpenStudy (anonymous):

f• g = f (g(x)) = 3 (7x+6) + 2 = 21x + 20

OpenStudy (anonymous):

\[\mathbb{R} = (-\infty, +\infty)\] , no restrictions for your function.